Two charges charge +q are kept d distence away. Mass of charges 2M and 3M. They are released from rest. Find time when separation becomes 3d.

Grade:12

1 Answers

Vikas TU
7 years ago
0 +  0 + q(kq/d) = 0.5*2M*v1^2 + 0.5*3M*v2^2 + q(kq/3d)
  0.5*2M*v1^2 + 0.5*3M*v2^2 = kq^2/d – kq^2/3d 
 0.5*M(2v1^2) + 0.5*M(3v2^2) = 2kq^2/3d ….......................(1)
 
Momentum conserved.
 
0  +  0 = 2M*v1  + 3M*v2
or
2v1 +  3v2 = 0 …..............(2)
 
solve eqns. (1) and (2)
and get v1 and v2.
 
calculate Time = 3d/v1or v2 (take any speed).